OpenStudy (anonymous):

Write the equation of the line in point-slope form that passes through (1, -4) and has a slope of 1/4

OpenStudy (anonymous):

point slope form: y-y1=m(x-x1) where (x1,y1) is the given point (1, -4) and m is the given slope of 1/4 (x1,y1) = (1, -4) m = 1/4 plug those values into y - y1 = m(x - x1)